OpenStudy (anonymous):

How do I rationalize 2 divided by 5 square root of 5

OpenStudy (anonymous):

\[\frac{ 2 }{ 5\sqrt{5} }\times \frac{ \sqrt{5} }{ \sqrt{5} }\]

OpenStudy (anonymous):

\[\frac{ 2\sqrt{5} }{ 25 }\]